OpenStudy (ttp):

A ball is thrown upward from the top of a building. The function below shows the height of the ball in relation to sea level, ft, in feet, at different times t in seconds: f(t) = -16t2 + 32t + 90 The average rate of change of f(t) from t = 4 seconds to t = 6 seconds is _____ feet per second.

OpenStudy (justin_lee):

Hey, just a question before I attempt to explain how to do this for you. Have you taken calculus or general physics? Do you know the average value formula?

OpenStudy (ttp):

No.

OpenStudy (justin_lee):

In that case, the rate of change is just the slope between the value of f(t) at 4 seconds and the value of f(t) at 6 seconds. Think of the seconds as the x-axis and f(t) as the y-axis. You know that a slope is: \[\frac{ y_{2} - y_{1} }{ x_{2} - x_{1} }\] In this case, your \[y_{2} = f(6 seconds)\] and your \[y_{1} = f(4 seconds)\] Your \[x_{2} = 6 seconds\] and your \[x_{1} = 4 seconds\] Does that help?